1. Harris Fabrics computes its predetermined overhead rate annually on the basis of direct labor-hours. At...

1.

Harris Fabrics computes its predetermined overhead rate annually on the basis of direct labor-hours. At the beginning of the year, it estimated that 37,000 direct labor-hours would be required for the period’s estimated level of production. The company also estimated $556,000 of fixed manufacturing overhead expenses for the coming period and variable manufacturing overhead of $3.00 per direct labor-hour. Harris's actual manufacturing overhead for the year was $737,889 and its actual total direct labor was 37,500 hours.


Required:

Compute the company's predetermined overhead rate for the year. (Round your answer to 2 decimal places.)

Predetermined overhead rate =    per DLH
Answer #1

Total variable overhead estimated=(3*37000)=$111,000

Total overhead estimated=Total variable overhead+Total fixed overhead

=111,000+556,000=$667,000

Predetermined overhead rate=Total overhead estimated/Total direct labor hours

=667,000/37,000

=$18.03 per DLH(Approx).